OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form.

OpenStudy (anonymous):

my options are y = 7x – 10 y = 7x + 10 y = 7x + 18 y = 7x – 18

OpenStudy (anonymous):

OH MY GOD ANYBODY CAN HELP ME ?

OpenStudy (anonymous):

y = 7x – 10 (2, 4) and (1, –3) 4= 7(2)-10 -3 =7(1) -10 The answer is what?

OpenStudy (anonymous):

The two points satisfy y = 7 x -10. The answer is y=7x-10